# Extreme softening of QCD phase transition under weak acceleration: first principle Monte Carlo results for gluon plasma

### v1

@@ Introduction.
The Wilson action (6) with two couplings ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ allows us to fix separately the spacial, ${\mathfrak{a}}_{\sigma}={\mathfrak{a}}_{\sigma}({\beta}_{\sigma},{\beta}_{\tau})$, and temporal, ${\mathfrak{a}}_{\tau}={\mathfrak{a}}_{\tau}({\beta}_{\sigma},{\beta}_{\tau})$, lattice spacings that stand for the physical length of the elementary lattice links in the appropriate directions.  [^2] To avoid confusion, we use the Fraktur font $\mathfrak{a}$ for a lattice spacing and the Italic font $a$ for an acceleration. We use an improved procedure following Ref. Karsch (1982) to simulate the theory on the asymmetric lattices. The technical details of our simulations are described in the Appendix.
We impose periodic boundary conditions along $x,\,y,\,\tau$ directions and keep open boundaries along the acceleration direction $z$. The physical length of the imaginary time circle, $L_{\tau}\equiv N_{\tau}{\mathfrak{a}}_{\tau}(z)=1/T(z)$ is chosen to match the temperature profile (4) corresponding to a constant proper acceleration. In the $z$ direction, temperature varies only on a central segment that occupies half of the volume. We keep two $\delta L_{z}={\mathfrak{a}}_{\sigma}N_{z}/4\sim(1\dots 4)\,{\rm fm}$ thick slices at both sites of the central segment to mitigate the effects of the open boundary conditions along the $z$ direction. It is important to stress that we made the lengths of the transverse spatial directions $L_{i}=N_{i}{\mathfrak{a}}_{\sigma}(z)$ with $i=x,y$ are made $z$-independent by imposing a particular $z$ dependence for the ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ couplings in the lattice action (6) (the details are given in the Appendix).
@@ Acknowledgements.
VAG and AVM have been supported by RSF (Project No. 23-12-00072, https://rscf.ru/project/23-12-00072/). DVS and ASP were supported by Grant No. FZNS-2024-0002 of the Ministry of Science and Higher Education of Russia. The numerical simulations were performed at the computing cluster of Far Eastern Federal University and the equipment of Shared Resource Center "Far Eastern Computing Resource" IACP FEB RAS (https://cc.dvo.ru).
Appendix
.1 Numerical simulations
The Wilson action (6) with two couplings ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ allows us to fix separately the spacial, ${\mathfrak{a}}_{\sigma}={\mathfrak{a}}_{\sigma}({\beta}_{\sigma},{\beta}_{\tau})$, and temporal, ${\mathfrak{a}}_{\tau}={\mathfrak{a}}_{\tau}({\beta}_{\sigma},{\beta}_{\tau})$ lattice spacings. The quantity
$\xi=\frac{{\mathfrak{a}}_{\sigma}}{{\mathfrak{a}}_{\tau}}\,,$ (10)
gives us the anisotropy of the lattice spacings.
In the continuum limit, where both couplings ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ are large, the lattice spacings vanish, ${\mathfrak{a}}_{\sigma}\to 0$ and ${\mathfrak{a}}_{\tau}\to 0$. Consequently, the lattice plaquette ${\mathcal{P}}_{x,\mu\nu}$ becomes proportional to the square of the continuum field-strength tensor $F_{\mu\nu}$ times a volume element, ${\mathfrak{a}}_{\sigma}^{3}{\mathfrak{a}}_{\tau}$, and the lattice Wilson action (6) reduces to the continuum action of the Yang-Mills theory Karsch (1982).
We work with asymmetric lattices, $N_{\tau}<N_{\sigma}$, corresponding to the case of finite temperature. The local temperature is given by an inverse of the lattice length $L_{\tau}=N_{\tau}{\mathfrak{a}}$ in the imaginary time direction:
$T(z)=\frac{1}{N_{\tau}{\mathfrak{a}}_{\tau}(z)}\,.$ (11)
We aim to introduce the temperature gradient along the longitudinal spatial direction $z\equiv x_{3}$:
$T(z)=\frac{T_{0}}{1+a_{0}(z-z_{0})}\,,$ (12)
where $T_{0}\equiv T(z{=}z_{0})$ is the temperature at the $z=z_{0}$ plane and $a_{0}>0$ is the parameter of the temperature inhomogeneity which corresponds, according to the Tolman-Ehrenfest relation (5), to the physical acceleration. We consider the case of a weak acceleration, $aL_{\sigma}\ll 1$, so that the physical temperature (12) is nowhere singular on our lattices. In other words, the position of the Rindler horizon, $z^{\rm R}=-1/a$, resides far away from the spatial region considered in our simulations, $|z^{\rm R}|\gg L_{\sigma}$.
We should choose the lattice couplings in Eq. (6) in such a way that the spatial lattice spacing remains constant along the whole lattice,
${\mathfrak{a}}_{\sigma}={\mathfrak{a}}_{0}=\frac{1}{N_{\tau}T_{0}}={\mathrm{const}}\,,$ (13)
while the physical temperature (11) varies according to Eq. (12), implying, in turn, that the temporal lattice spacing is a linear function of the longitudinal coordinate $z$:
${\mathfrak{a}}_{\tau}(z)={\mathfrak{a}}_{0}\bigl{(}1+a_{0}(z-z_{0})\bigr{)}\,,$ (14)
which then corresponds to the anisotropy parameter (10):
$\xi(z)=\frac{1}{1+a_{0}(z-z_{0})}=\frac{T(z)}{T_{0}}\,.$ (15)
We work at fixed lattice geometry $N_{\tau}N_{\sigma}^{3}$ so that the variation of temperature in the $z$ direction is assumed only in the lattice couplings ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ while the lattice extension in the temporal direction, $N_{\tau}$, is obviously independent of $z$. For the sake of convenience, we take $z_{0}=0$ in Eqs. (12), (14), and (15).
A temperature gradient has also been implemented in a recent study of another problem Yang et al. (2024) without imposing the fixed-spatial constraint (13) potentially generating unwanted varying-volume effects. In addition, we notice that the non-linear trigonometric behavior of temperature inhomogeneities introduced in Ref. Yang et al. (2024) differs from the thermodynamic formula (4) given by the solution of the Killing equation (2). Therefore, the trigonometric temperature inhomogeneities studied in Ref. Yang et al. (2024) do not correspond to the global thermal equilibrium of the accelerating medium studied in our article.
.2 Setting up the couplings
One can consider a linear approximation with Eqs. (K.2.4), (K.2.24), and (K.2.25) of Ref. Karsch (1982), where “K” in front of the equation corresponds to this paper by F. Karsch. It is convenient to parameterize the couplings as follows Karsch (1982):
${\beta}_{\sigma}=\frac{6}{g_{\sigma}^{2}({\mathfrak{a}},\xi)}\frac{1}{\xi}\,,\qquad{\beta}_{\tau}=\frac{6}{g_{\tau}^{2}({\mathfrak{a}},\xi)}\xi\,,$ (16)
where at the symmetric value, $\xi=1$, the couplings are equal:
$g_{\tau}({\mathfrak{a}},1)=g_{\sigma}({\mathfrak{a}},1)=g_{E}({\mathfrak{a}})\,.$ (17)
For an asymmetric lattice, $\xi\neq 1$, the lattice couplings ${\beta}_{\tau}$ and $\beta_{\sigma}$ become different from each other. In the weak coupling limit, relevant to the continuum limit, they can be expanded in terms of the single symmetric coupling $g_{E}$:
$\frac{1}{g_{\sigma}^{2}({\mathfrak{a}},\xi)}$ $=\frac{1}{g_{E}^{2}({\mathfrak{a}})}+c_{\sigma}(\xi)+O(g_{E}^{3})\,,$ (18a)
$\frac{1}{g_{\tau}^{2}({\mathfrak{a}},\xi)}$ $=\frac{1}{g_{E}^{2}({\mathfrak{a}})}+c_{\tau}(\xi)+O(g_{E}^{3})\,,$ (18b)
where the functions $c_{\sigma}(\xi)$ and $c_{\tau}(\xi)$ are given in Ref. Karsch (1982). Thus, one should collect all relevant equations of Ref. Karsch (1982), neglect $O(g_{E}^{3})$ terms in Eqs. (18) and use the anisotropy parameter (15) to generate the temperature gradient (12). However, one should do this carefully.
We rewrite Eqs. (16) and (18) as follows:
${\beta}_{\sigma}(\beta,\xi)$ $=\frac{\beta}{\xi}+\frac{6}{\xi}c_{\sigma}(\xi)\,,$ (19a)
${\beta}_{\tau}(\beta,\xi)$ $=\beta\xi+6\xi c_{\tau}(\xi)\,,$ (19b)
where $O(g_{E}^{3})$ terms are neglected. We also denoted
${\beta}_{\sigma}=\frac{6}{g_{\sigma}^{2}}\frac{1}{\xi}\,,\qquad{\beta}_{\tau}=\frac{6}{g_{\tau}^{2}}\xi\,,\qquad\beta=\frac{6}{g_{E}^{2}}\,,$ (20)
where $\beta$ plays a role of the central value of the lattice coupling according to Eqs. (19). From now on, we will work only in terms of $\beta$ and $\xi$.
We must work with the following conditions:
  1. (i)
    one has anisotropy between spatial and temporal lattice spacings (10);
  2. (ii)
    the spatial lattice spacing is fixed regardless of the value of the anisotropy.
In other words, we have to find the set of the lattice couplings $\beta$ and $\xi$ which fulfill the following conditions:
  1. (i)
    The anisotropy (10) is given by the value of $\xi$:
    $\xi=\frac{{\mathfrak{a}}_{\sigma}(\beta,\xi)}{{\mathfrak{a}}_{\tau}(\beta,\xi)}\,.$ (21)
  2. (ii)
    At a fixed $\xi$, the value of $\beta$ should be chosen in such a way that the spatial lattice spacing (13) is independent of the asymmetry $\xi$ so that the variations in temperature do not affect the spatial geometry (size of the spatial directions) of the system:
    ${\mathfrak{a}}_{\sigma}(\beta,\xi)={\mathfrak{a}}_{0}\,.$ (22)
Condition (i) is already ensured by the choice of Eq. (19). However, condition (ii) is less trivial and has to be explicitly cared for.
In the case of the absence of anisotropy, $\xi=1$, Eq. (22) means that a fixed value of ${\mathfrak{a}}$ is achieved at a fixed value of $\beta$. In the presence of a nontrivial anisotropy, $\xi\neq 1$, Eq. (22) implies that a fixed ${\mathfrak{a}}$ constraints a relation $\beta={\beta}_{f}(\xi,{\mathfrak{a}}_{0})$ between $\beta$ and $\xi$, such that ${\mathfrak{a}}_{\sigma}(\beta(\xi,{\mathfrak{a}}_{0}),\xi)={\mathfrak{a}}_{0}$. Here the subscript “$f$” in ${\beta}_{f}$ means “fixed”.
Coming back to Ref. Karsch (1982), we notice from Eq. (K.2.6) that the spatial lattice spacing has the following behavior:
${\mathfrak{a}}(\beta,\xi)={\mathfrak{a}}(\beta)\exp\left[-\frac{c_{\sigma}(\xi)+c_{\tau}(\xi)}{4b_{0}}\right]\,,$ (23)
where ${\mathfrak{a}}(\beta)$ is the lattice spacing for the isotropic lattice, ${\mathfrak{a}}(\beta)\equiv{\mathfrak{a}}(\beta,\xi=1)$ and
$b_{0}=\frac{11N_{c}}{48\pi^{2}}\equiv\frac{11}{16\pi^{2}}=0.69\dots\qquad\text{for}\quad N_{c}=3\,,$ (24)
is the first coefficient of the beta function. According to Ref. Karsch (1982), the correction in the spatial lattice spacing (23) due to non-trivial behavior of $\xi$ is very small, of the order of 1%. The physical values of the isotropic lattice coupling ${\mathfrak{a}}(\beta)$ in Eq. (23) as functions of $\beta$ can be found in Table 1 of Ref. Athenodorou and Teper (2020).
The final functions defining the $\beta(\xi)$ dependence are presented below:
${\beta}_{\sigma}({\beta}_{0},\xi)$ $=\frac{1}{\xi}\left(\beta({\beta}_{0},\xi)+6\cdot 1.105\cdot c_{\sigma}(\xi)\right)\,,$ (25a)
${\beta}_{\tau}({\beta}_{0},\xi)$ $=\xi\left(\beta({\beta}_{0},\xi)+6\cdot c_{\tau}(\xi)\right)\,,$ (25b)
$\beta({\beta}_{0},\xi)$ $=\beta\left({\mathfrak{a}}({\beta}_{0})\exp\left[\frac{c_{\sigma}(\xi)+c_{\tau}(\xi)}{4b_{0}}\right]\right)\,,$ (25c)
where function $\beta\left({\mathfrak{a}}\right)$ is the inverse function for ${\mathfrak{a}}\left(\beta\right)$ Athenodorou and Teper (2020), ${\beta}_{0}$ is the critical $\beta$-value for finite temperature. We performed calculations with the lattice size in temporal direction $N_{t}=8$ (${\beta}_{0}=6.07$). In function (25a) we add the additional multiplication factor $1.105$ for improve stability of the ${\mathfrak{a}}_{\sigma}={\mathrm{const}}$ in the used $\xi$-range. We got the value of the factor by fitting lattice data.
At zero temperature, we explicitly calculated the asymmetry value $\xi$ from the Creutz ratio shown in Fig. 5. We kept the spatial size of the lattice constant, and the resulting deviations were, again, less than one percent.
Figure 5: Relations ${\mathfrak{a}}_{\sigma}/{\mathfrak{a}}_{\tau}$ and ${\mathfrak{a}}_{\sigma}/{\mathfrak{a}}_{0}$ are obtained from the analysis of the Creutz ratio for Wilson loops $5\times 5$ for ${\beta}_{0}=6.07$ on the lattice $42^{4}$. We used the line function for fit: $y(\xi)=k\xi+(1-k)$ which gives us the following best fit value: $k=0.600(16)$.

### v2

@@ Introduction.
The Wilson action (6) with two couplings ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ allows us to fix separately the spacial, ${\mathfrak{a}}_{\sigma}={\mathfrak{a}}_{\sigma}({\beta}_{\sigma},{\beta}_{\tau})$, and temporal, ${\mathfrak{a}}_{\tau}={\mathfrak{a}}_{\tau}({\beta}_{\sigma},{\beta}_{\tau})$, lattice spacings that stand for the physical length of the elementary lattice links in the appropriate directions.  [^2] To avoid confusion, we use the Fraktur font $\mathfrak{a}$ for a lattice spacing and the Italic font $a$ for an acceleration. We use an improved procedure following Ref. Karsch (1982) to simulate the theory on the asymmetric lattices. The technical details of our simulations are described in the Supplemental Material.
We impose periodic boundary conditions along $x,\,y,\,\tau$ directions and keep open boundaries along the acceleration direction $z$. The physical length of the imaginary time circle, $L_{\tau}\equiv N_{\tau}{\mathfrak{a}}_{\tau}(z)=1/T(z)$ is chosen to match the temperature profile (4) corresponding to a constant proper acceleration. In the $z$ direction, temperature varies only on a central segment that occupies half of the volume. We keep two $\delta L_{z}={\mathfrak{a}}_{\sigma}N_{z}/4\sim(1\dots 4)\,{\rm fm}$ thick slices at both sites of the central segment to mitigate the effects of the open boundary conditions along the $z$ direction. It is important to stress that we made the lengths of the transverse spatial directions $L_{i}=N_{i}{\mathfrak{a}}_{\sigma}(z)$ with $i=x,y$ are made $z$-independent by imposing a particular $z$ dependence for the ${\beta}_{\sigma}$ and ${\beta}_{\tau}$ couplings in the lattice action (6) (the details are given in the Supplemental Material).
@@ Acknowledgements.
The work of MNC has been supported by the European Union - NextGenerationEU through grant No. 760079/23.05.2023, funded by the Romanian ministry of research, innovation and digitalization through Romania’s National Recovery and Resilience Plan, call no. PNRR-III-C9-2022-I8. VAG and AVM have been supported by RSF (Project No. 23-12-00072, https://rscf.ru/project/23-12-00072/). DVS and ASP were supported by Grant No. FZNS-2024-0002 of the Ministry of Science and Higher Education of Russia. The numerical simulations were performed at the computing cluster of Far Eastern Federal University and the equipment of Shared Resource Center "Far Eastern Computing Resource" IACP FEB RAS (https://cc.dvo.ru).